A 100 W bulb has tungsten filament of total length 1.0 m and radius 4× 10-5 m. The emissivity of the filament is 0.8 and σ = 6.0 × 10-8 W/m2-K4. Calculate the temperature of the filament when the bulb is operating at correct wattage.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Q/t = Ae σT^4 ⇒ T^4 = θ/teAσ ⇒ T^4 = 100/0.8 * 2 *3.14 * 4 *10^-5 * 1 * 6 * 10^-8 ⇒ T = 1697.0 = 1700 K